Solve the equation:
5y - 10 = -9 - 4y
New to these sort of problems. Thanks for your time.

Found 2 solutions by Alan3354, lenny460:
Answer by Alan3354(69443)

Add 4y to each side.
Then add 10 to each side.

Answer by lenny460(1073)

Solve the equation:
5y - 10 = -9 - 4y
New to these sort of problems. Thanks for your time.
Add 4y to each side of the equation.
5y - 10 + 4y = -9 - 4y + 4y
9y - 10 = -9
Add 10 to each side of the equation.
9y - 10 + 10 = -9 + 10
9y = 1
Divide each side by 9.
9y/9 = 1/9


Answer:
y = 1/9